A country produces only bananas and coconuts. The base year is 2015, and the table below gives the quantities produced and prices.

 

 

Year

Bananas bunches

Coconuts bunches

Quantities produced

2015

1,000

500

Quantities produced

2016

1,100

525

Prices

2015

$2

$10

Prices

2016

$3

$8

 

  1. Calculate nominal GDP in 2015 and 2016:
  2. Calculate the RGDP 2016 
1
Expert's answer
2021-04-06T07:27:07-0400

1) Let's calculate the nominal GDP in 2015 and 2016:


"Nominal\ GDP_{2015}=1000\ b \cdot\dfrac{\$2}{b}+500\ c\cdot\dfrac{\$10}{c}=\$7000,""Nominal\ GDP_{2016}=1100\ b \cdot\dfrac{\$3}{b}+525\ c\cdot\dfrac{\$8}{c}=\$7500."

2) Let's calculate the real GDP in 2016:


"Real\ GDP_{2016}=1100\ b \cdot\dfrac{\$2}{b}+525\ c\cdot\dfrac{\$10}{c}=\$7450."